OpenStudy (shaleiah):

At 7:00 pm the temperature is 72⁰F. At 10:00 pm it is 52⁰F. Find the average rate of change in the temperature. @jmthompson5910

OpenStudy (shaleiah):

@jim_thompson5910

OpenStudy (shaleiah):

@IrishBoy123

OpenStudy (irishboy123):

ideas?!

OpenStudy (irishboy123):

\(slope = \dfrac{rise}{run}\)

OpenStudy (shaleiah):

20/4=5 ?

OpenStudy (irishboy123):

\( \dfrac{rise}{run} = \dfrac{52 - 72}{10 - 7}\)

OpenStudy (shaleiah):

-6.67

OpenStudy (irishboy123):

yes! units??

OpenStudy (shaleiah):

Fahrenheit

OpenStudy (irishboy123):

per hour

OpenStudy (shaleiah):

Thank you

OpenStudy (irishboy123):

np